# LLM-Assisted Request Mutation for Whitebox REST API Testing

### v2

@@ Abstract.
Cloud applications heavily rely on APIs to communicate with each other and exchange data. To ensure the reliability of cloud applications, cloud providers widely adopt API testing techniques. Unfortunately, existing API testing approaches are insufficient to reach strict conditions, a problem known as fitness plateaus, due to the lack of gradient provided by coverage metrics. To address this issue, we propose MioHint, a novel white-box API testing approach that leverages the code comprehension capabilities of Large Language Model (LLM) to boost API testing. The key challenge of LLM-based API testing lies in system-level testing, which emphasizes the dependencies between requests and targets across functions and files, thereby making the entire codebase the object of analysis. However, feeding the entire codebase to an LLM is impractical due to its limited context length and short memory. MioHint addresses this challenge by synergizing static analysis with LLMs. We retrieve relevant code with data-dependency analysis at the statement level, including def-use analysis for variables used in the target and function expansion for subfunctions called by the target.
To evaluate the effectiveness of our method, we conducted experiments across 16 real-world REST API services. The findings reveal that MioHint achieves an average increase of 4.95% absolute in line coverage compared to the baseline, EvoMaster, alongside a remarkable factor of 67$\times$ improvement in mutation accuracy. Furthermore, our method successfully covers over 57% of hard-to-cover targets while in baseline the coverage is less than 10%.
@@ 1. Introduction
Despite its importance, the majority of automated testing efforts are still centered around black-box testing (Atlidakis et al., 2019; Liu et al., 2022; Viglianisi et al., 2020; Golmohammadi et al., 2022), which typically achieves low coverage (Zhang and Arcuri, 2022). Black-box API testing treats the API as a closed system, focusing solely on the inputs and outputs without any knowledge of the internal workings of the API. While this approach is valuable for validating the external behavior of the API, it often falls short in dealing with corner cases and deep system states.
On the other hand, white-box API testing leverages runtime information (e.g., coverage) of the APIs by instrumentation of source code. With this information, white-box testing defines heuristics and applies search-based techniques to fuzzing REST APIs. This approach leads to significantly higher results in code coverage because of the search guidance provided by coverage metrics (Zhang and Arcuri, 2022). However, white-box approaches fall short when facing fitness plateaus (Aleti and Grunske, 2017), a widely recognized problem that traps the input search process in local optima, where coverage provides no gradient to the search, e.g., a strict condition like checking if the input equals a specific value.
To address these limitations, there is a growing need for more sophisticated white-box testing approaches that leverage a deeper understanding of the codebase. One such approach is symbolic execution, which tracks program inputs as symbolic variables and maintains symbolic expressions across statements. This technique allows for the construction of constraints on the input related to the target, which can then be solved using a constraint solver.
@@
To address the common issue of fitness plateaus in search algorithms, we enhance search accuracy by a lightweight analysis provided by the LLMs. Also, we retrieve global context through statement-level data dependency analysis to effectively capture the relevant context while minimizing redundancy.
Our proposed approach, MioHint, initiates LLM-assisted mutation when the search algorithm encounters hard-to-cover targets. For each target, MioHint extracts the relevant code by employing statement-level value expansion to generate global context. It then constructs a prompt using this context and queries GPT-4o for mutation guidance. This approach allows us to utilize the search algorithm’s fitness function to evaluate the quality of tests generated by GPT-4o, followed by additional random mutations on the high-quality test cases produced by the LLM.
We integrate MioHint into EvoMaster (Arcuri et al., 2021), a widely adopted whitebox API testing framework. We evaluate it with 16 real-world REST API services of EMB (EMB, [n. d.]), which add up to 314,415 lines of code. The experiments demonstrate that compared to the baseline EvoMaster, our approach increases line coverage by an average of 4.92% and achieves a 67$\times$ increase in mutation accuracy. Besides, our approach successfully covers over 57% of hard-to-cover targets while in baseline the coverage is less than 10%.
@@
  • We propose MioHint which integrates LLM-assisted mutation with an advanced search algorithm. It makes a substantial improvement on the accuracy of mutation with the aid of LLMs.
@@ 2.1. Web Services and API
Web services and APIs (Application Programming Interfaces) are essential components in modern software development. As shown in Figure  1, APIs provide standardized methods for applications to exchange data and functionality, regardless of the underlying platforms or technologies.
The REST API (Representational State Transfer Application Programming Interface) is the most widely adopted API specification. REST API is based on the HTTP protocol. It typically employs standard HTTP methods (GET, POST, PUT, DELETE) to define operations, uses URIs to identify resources, leverages standard HTTP status codes (such as 200, 404, 500) to indicate request outcomes, utilizes JSON or XML for data interchange, and incorporates HTTP headers (like Content-Type and Authorization) to convey metadata and control information. We listed examples of request and response in REST APIs in Figure  1, to show how applications exchange data by REST APIs.
@@ 2.2. Automated API Testing
BlackBox API Testing. Most of the existing API testing (i.e., 73%) are black-box (Golmohammadi et al., 2022). BlackBox API testing conceptualizes the SUTs as an enclosed entity. It validates the system’s behavior by iteratively initiating API requests and verifying the correctness of their responses.
The key challenges for black-box approaches are 1) generating valid inputs and 2) defining appropriate test oracles to evaluate system correctness. Existing approaches rely on API specifications provided by service providers to generate valid inputs. They employ machine learning techniques or regular expression matching to interpret the specifications and subsequently generate valid requests (Liu et al., 2022; Atlidakis et al., 2019). Regarding test oracles, they typically depend on the status codes embedded in API responses to check the system’s correctness, e.g., 500 refers to an internal server error.
Unfortunately, existing black-box testing approaches suffer from low code coverage since they are invisible to the system implementations. Consequently, they can only randomly mutate inputs based on limited information in specifications. This leaves complex system states surrounded by strict constraints and corner cases uncovered. Additionally, these approaches heavily depend on accurate and complete specifications from service providers, a condition that is often unmet. When faced with inaccurate specifications, existing methods tend to generate numerous invalid inputs, further diminishing code coverage (Zhang and Arcuri, 2022; Martin-Lopez et al., 2021). Besides, the under-specified schemas result in all the code related to handling and associated functionalities never being tested if certain HTTP headers and query parameters are not defined in the specification (Arcuri et al., 2023).
WhiteBox API Testing. White-box approaches achieve better code coverage by inspecting the runtime information of SUTs. EvoMaster is the state-of-the-art white-box API fuzzer that adopts code instrumentation to increase the internal visibility of SUTs (Arcuri et al., 2021). Specifically, EvoMaster first identifies the code not yet covered by maintaining an access counter at each line of bytecode. The counter is updated each time the bytecode is executed. It then adopts a heuristic-based search algorithm named Many Independent Object (MIO) to maximize code coverage with the collected bytecode-level coverage information.
@@
Figure 2. High-level view of Search-based API Testing.
Specifically, MIO is a variation of the evolutionary algorithm. It maintains a pool of high-quality inputs, i.e., inputs that have the potential to increase the coverage. Each time during testing, MIO randomly selects an input from the pool, mutates the input to let it execute different parts of code, and examines the code coverage during execution. A mutated input is added to the input pool only if it increases code coverage. This is because inputs that increase code coverage imply that they have triggered the execution of code blocks unseen before.
Unfortunately, the code coverage of EvoMaster based on MIO is still unsatisfactory. The fundamental reason is due to the ubiquitous fitness plateaus, i.e., mutated test cases are unlikely to increase coverage, and the search process is trapped in local optima (Aleti and Grunske, 2017).
@@
A straightforward approach is to adopt traditional code analysis techniques, e.g., symbolic execution, to overcome the fitness plateaus. Symbolic execution tracks the semantics in terms of any symbolic input at the bytecode level, formulates constraints, and satisfies them by an SMT solver (De Moura and Bjørner, 2008). However, the application of symbolic execution to complex real-world software inevitably encounters several challenges, including model boundaries (e.g., external libraries), path explosion, and imprecise abstraction (Yun et al., 2018; Poeplau and Francillon, 2020). Specifically, when external libraries are not analyzed, the symbolic execution loses track of symbolic expressions whenever data is passed through a function provided by the library. Besides, the number of control-flow paths grows exponentially with an increase in program size. Moreover, any imprecise abstraction of complex data types can lead to incorrect constraints.
@@ 2.3. Opportunity of code understanding with LLM
We propose to use the code comprehension capabilities of LLMs to boost the coverage of API testing. An LLM is an auto-regressive model designed to understand and generate natural language by leveraging vast amounts of data (e.g., 700 GB), it presents an impressive performance at code understanding and generation (Liu et al., 2024a; Qiu et al., 2024).
Using LLMs can effectively address the issue of fitness plateaus without the additional overhead of symbolic execution. First, an LLM analyzes static code snippets instead of resolving all data types and dependencies at the bytecode level, thus reducing the analysis overhead. Second, LLM can directly extract the functionality of external function calls by their name matches, parseInt, and group instead of symbolically reasoning about them.
As shown in Figure 3, we use LLM to generate input for the same example. We query the LLM for mutation hints with function code, target line code, and an initial input to mutate. LLM effectively generates a hint in the label (2), which reaches cPos < 1 == true in a single query. LLMs achieve this by understanding the functionality of our study function from a semantic perspective and directly generating a string c.0A>G that fits the requirement.
Finally, a new test case is generated by applying the mutation hint to the selected API call, labeled (3). Once the high-quality mutant is generated, it is immediately evaluated to determine any increase in fitness. If the fitness improves, the high-quality mutant is saved as a new test case. Subsequently, random mutations can be performed on this new test case to explore other parts of the search space. This approach effectively escapes local optima and yields additional benefits beyond merely covering one target.
@@ 2.4. Challenges of LLM-Based API Testing at Repository Level
Large language models have been widely adopted to unit test generation for their remarkable capabilities in code understanding and generation. To enhance the effectiveness of unit test generation, numerous methods have been proposed (Alagarsamy et al., 2023; Shin et al., 2024; Khanfir et al., 2023; Chen et al., 2024; Yuan et al., 2024; Vikram et al., 2024; Plein et al., 2023; Lemieux et al., 2023). However, unit testing primarily focuses on isolated functions and small segments of code within a repository and has no concern about other parts of the codebase. In contrast, our API testing is a system-level testing, that emphasizes the dependencies between requests and targets. These dependencies span across functions and files, thereby constituting a repository-level concern.
Repository-level tasks are challenging as information spans across massive code. State-of-the-art works, including package migration  (Bairi et al., 2023), issue resolution  (Ma et al., 2024; Xia et al., 2024; Zhang et al., 2024), and code completion  (Phan et al., 2024; Ding et al., 2023; Shrivastava et al., 2023), employs dependency graphs to identify relevant code components through method/class relationships (e.g., imports, invocations, inheritance).
However, different from existing repository-level tasks that focus on capturing module interactions, API testing emphasizes value propagation paths between API requests and testing targets. Specifically, the goal of API testing is to cover the target by modifying the request. Therefore, the key is to determine the target’s expression with respect to the request. Once this expression is established, it provides a constraint representing the target condition. To cover the target, we simply need to satisfy this constraint. To build the expression with respect to the request, we need to track value transmission at the statement level. This analysis works as starting from the variables of the target and iteratively finds their definitions until the request variable. Previous method or class level dependency graphs, though useful for identifying related modules, introduce extraneous code elements that obscure the critical value transmission chain. This over-inclusion occurs because class/method dependencies aggregate all interactions within a component, whereas only a subset of these interactions pertain to the specific request-target relationship.
Considering the accuracy of semantic representation, as well as the limitations on the length of LLM context and cost considerations, we have opted to conduct code extraction at the statement level.
@@ 3. MioHint
To improve mutation accuracy, we substitute vanilla mutation with our LLM-assisted mutation. LLM-assisted mutation comprises three main steps: related code extraction, prompt construction, and test case generation. Related code extraction retrieves static code related to the target statement from the codebase. Prompt construction combines the information gathered to form a task instruction. Test generation queries LLMs with the constructed prompt, processes the response to get the mutation hint, and applies the mutation hint to the original test case to generate a new test case.
In the rest of this section, we first elaborate MioHint at a high level, and then focus on the related code extraction, prompt construction, and test generation in LLM-assisted mutation.

### v3

# MioHint: LLM-Assisted Request Mutation for Whitebox REST API Testing

@@ Abstract.
Cloud applications heavily rely on APIs to communicate with each other and exchange data. To ensure the reliability of cloud applications, cloud providers widely adopt API testing techniques. Unfortunately, existing API testing approaches are insufficient to reach strict conditions, a problem known as fitness plateaus, due to the lack of a gradient provided by coverage metrics. To address this issue, we propose MioHint, a novel white-box API testing approach that leverages the code comprehension capabilities of Large Language Model (LLM) to boost API testing. The key challenge of LLM-based API testing lies in system-level testing, which emphasizes the dependencies between requests and targets across functions and files, thereby making the entire codebase the object of analysis. However, feeding the entire codebase to an LLM is impractical due to its limited context length and short memory. MioHint addresses this challenge by synergizing static analysis with LLMs. We retrieve relevant code with data-dependency analysis at the statement level, including def-use analysis for variables used in the target and function expansion for subfunctions called by the target.
[…]
API Testing, Large Language Models, Code Comprehension, Static Analysis
$c$s: Software and its engineering Software testing and debugging
@@ 1. Introduction
Despite its importance, the majority of automated testing efforts are still centered around black-box testing (Atlidakis et al., 2019; Liu et al., 2022; Viglianisi et al., 2020; Golmohammadi et al., 2023), which typically achieves low coverage (Zhang and Arcuri, 2022). Black-box API testing treats the API as a closed system, focusing solely on the inputs and outputs without any knowledge of the internal workings of the API. While this approach is valuable for validating the external behavior of the API, it often falls short in dealing with corner cases and deep system states.
On the other hand, white-box API testing leverages runtime information (e.g., coverage) of the APIs by instrumentation of source code. With this information, white-box testing defines heuristics and applies search-based techniques to fuzz REST APIs. Specifically, it generates new test cases by randomly mutating existing test cases, and then evaluates the new test cases with its code coverage. This approach achieves better results in code coverage due to the coverage-guided search process (Zhang and Arcuri, 2022). However, white-box approaches fall short when encountering fitness plateaus (Arcuri and Galeotti, 2021), a widely recognized problem when coverage provides no gradient to the search, e.g., a strict equality check. In this situation, random mutations become inefficient at improving coverage.
To address these limitations, there is a growing need for more sophisticated white-box testing approaches that leverage a deeper understanding of the codebase. One such approach is symbolic execution, which tracks program inputs as symbolic variables and maintains symbolic expressions across statements. This technique allows for the construction of constraints on the input related to the target, which can then be solved using a constraint solver. After solving, the program can directly reach the target with the solution.
@@
To address the common issue of fitness plateaus in search algorithms, we enhance mutation accuracy with a lightweight analysis provided by the LLMs. Ideally, high-accuracy mutations can directly reach hard-to-cover targets, thereby overcoming the entrapment in local optima imposed by fitness plateaus. Given the vast information landscape presented by system-level testing, specifically the entire repository, we retrieve global context through statement-level data dependency analysis to effectively capture the relevant context while minimizing redundancy.
Our proposed approach, MioHint, initiates LLM-assisted mutation when the search algorithm encounters hard-to-cover targets. This integrated approach creates a synergistic system that balances search effectiveness with efficiency. It leverages the fast, broad exploration capabilities of the search algorithm for general progress, while strategically deploying the targeted LLM-assisted mutation as a powerful specialist to overcome difficult coverage targets that would otherwise cause the search to stagnate. This division of labor effectively optimizes the trade-off between search breadth and depth, thereby maximizing overall testing performance. In the implementation of LLM-assisted mutation, for each target, MioHint begins by performing statement-level value expansion to extract global context from the entire codebase. It then constructs a sophisticated prompt by synthesizing this global context with local function code and existing requests. This prompt strategically employs chain-of-thought reasoning to guide the LLM in methodically locating the field for revision and in-context learning to incorporate feedback from previous failed mutations. Finally, this structured prompt is used to query the LLM for precise mutation guidance.
[…]
@@
  • We propose MioHint, which integrates LLM-assisted mutation with an advanced search algorithm. It makes a substantial improvement on the accuracy of mutation with the aid of LLMs.
@@ 2.1. Web Services and API
Web services and APIs (Application Programming Interfaces) are essential components in modern software development. As shown in Figure 1, APIs provide standardized methods for applications to exchange data and functionality, regardless of the underlying platforms or technologies.
The REST API (Representational State Transfer Application Programming Interface) is the most widely adopted API specification. REST API is based on the HTTP protocol. It typically employs standard HTTP methods (GET, POST, PUT, DELETE) to define operations, uses URIs to identify resources, leverages standard HTTP status codes (such as 200, 404, 500) to indicate request outcomes, utilizes JSON or XML for data interchange, and incorporates HTTP headers (like Content-Type and Authorization) to convey metadata and control information. We listed examples of request and response in REST APIs in Figure 1, to show how applications exchange data by REST APIs.
Figure 2. High-level view of Search-based API Testing.
@@ 2.2. Automated API Testing
BlackBox API Testing. Most of the existing API testing tools (i.e., 73%) are black-box (Golmohammadi et al., 2023). BlackBox API testing conceptualizes the SUTs as an enclosed entity. It validates the system’s behavior by iteratively initiating API requests and verifying the correctness of their responses. The key challenges for black-box approaches are 1) generating valid inputs and 2) defining appropriate test oracles to evaluate system correctness. Existing approaches rely on API specifications provided by service providers to generate valid inputs. They employ machine learning techniques or regular expression matching to interpret the specifications and subsequently generate valid requests (Liu et al., 2022; Atlidakis et al., 2019). Regarding test oracles, they typically depend on the status codes embedded in API responses to check the system’s correctness, e.g., 500 refers to an internal server error.
Existing black-box testing approaches often achieve low code coverage because they lack access to the system’s implementation. Consequently, they can only mutate inputs randomly based on the limited information provided in specifications. This random process frequently fails to cover complex system states, strict constraints, and corner cases. Furthermore, these approaches rely heavily on accurate and complete specifications from service providers, a requirement that is often not met. When specifications are inaccurate, these methods tend to generate numerous invalid inputs, which further diminishes code coverage (Zhang and Arcuri, 2022; Martin-Lopez et al., 2021). Moreover, the under-specified schema results in all the code related to handling and associated functionalities never being tested if certain HTTP headers and query parameters are not defined in the specification (Arcuri et al., 2023).
WhiteBox API Testing. White-box approaches achieve higher code coverage by inspecting the runtime information of SUTs. For instance, EvoMaster is the state-of-the-art white-box API fuzzer that adopts code instrumentation to gain the internal visibility into the SUTs (Arcuri et al., 2021). Specifically, EvoMaster identifies uncovered code by maintaining an access counter for each line of bytecode. This counter is incremented each time the corresponding bytecode is executed. Subsequently, it employs a heuristic-based search algorithm, Many Independent Objects (MIO), which uses the collected bytecode-level coverage information to maximize code coverage.
@@
Specifically, MIO is a variation of the evolutionary algorithm. It maintains a pool of high-quality inputs, i.e., inputs that have the potential to increase the coverage. In each iteration, the algorithm randomly selects an input from this pool and applies mutations to it. The goal of this mutation is to execute different parts of the code. The code coverage is then measured during its execution. A mutated input is added to the pool only if it increases the overall code coverage. The rationale behind this is that such inputs have successfully triggered previously unexecuted code blocks.
However, the code coverage achieved by EvoMaster using MIO is still unsatisfactory. The fundamental reason is due to the ubiquitous fitness plateaus, i.e., mutated test cases are unlikely to increase coverage, and the search process is trapped in local optima (Aleti and Grunske, 2017).
@@
A straightforward approach is to adopt traditional code analysis techniques, e.g., symbolic execution, to overcome the fitness plateaus. Symbolic execution tracks the semantics in terms of any symbolic input at the bytecode level, formulates constraints, and satisfies them by an SMT solver (De Moura and Bjørner, 2008). However, the application of symbolic execution to complex real-world software inevitably encounters several challenges, including model boundaries (e.g., external libraries), path explosion, and imprecise abstraction (Yun et al., 2018; Poeplau and Francillon, 2020). Specifically, when external libraries are not analyzed, the symbolic execution loses track of symbolic expressions whenever data is passed through a function provided by the library. Besides, the number of control-flow paths grows exponentially with an increase in program size, resulting in significant analysis overhead. Moreover, any imprecise abstraction of complex data types can lead to incorrect constraints.
@@ 2.3. Opportunity of Code Understanding with LLM
We propose leveraging the code comprehension capabilities of LLMs to boost the coverage of API testing. LLMs are auto-regressive models designed to understand and generate natural language by leveraging vast amounts of data (e.g., 700 GB). It presents an impressive performance at code understanding and generation (Liu et al., 2024a; Qiu et al., 2024).
Using LLMs can effectively address the issue of fitness plateaus without the additional overhead of traditional techniques like symbolic execution. First, an LLM analyzes fragmented code snippets instead of resolving all data types and dependencies at the bytecode level, thus reducing the analysis overhead. Second, LLMs can infer the functionality of external function calls from their names, such as matches, parseInt, and group, instead of performing complex symbolic reasoning on them.
As shown in Figure 3, we leverage the LLM to generate input for the same example. The LLM is queried with the function code, the target line of code, and an initial input to mutate. Remarkably, the LLM produces the mutation hint shown at label (2), enabling the search to reach the condition cPos < 1 == true in a single query. This is achieved through the LLM’s semantic understanding of the function: it directly generates the string c.0A>G for field hgvsc, which satisfies the target condition.
Finally, a new test case is generated by applying the mutation hint to the selected API call, labeled (3). This new test case is crucial for breaking through the fitness plateau and improving code coverage. This is because it successfully exercises a branch of the code that was previously unreachable by random mutations.
The LLM-generated hint is effective for two reasons. First, by analyzing data dependencies in the code, the LLM correctly identifies that the input field hgvsc influences the critical variable cPos in the target condition. Subsequently, it determines the precise value c.0A>G required to steer program execution to a previously uncovered branch by satisfying the condition cPos < 1.
Once the high-quality mutant is generated, it is immediately evaluated to determine any increase in fitness. If the fitness improves, the high-quality mutant is saved as a new test case. Subsequently, random mutations can be performed on this new test case to explore other parts of the search space broadly. This approach effectively escapes local optima and yields additional benefits beyond merely covering one target.
@@ 2.4. Challenges of LLM-Based API Testing at Repository Level
Large language models have been widely adopted for unit test generation due to their remarkable capabilities in code understanding and generation. To enhance the effectiveness of unit test generation, numerous methods have been proposed (Alagarsamy et al., 2024; Shin et al., 2024; Khanfir et al., 2023; Chen et al., 2024; Yuan et al., 2024; Vikram et al., 2024; Plein et al., 2023; Lemieux et al., 2023). However, unit testing primarily focuses on isolated functions and small segments of code within a repository and has no concern about other parts of the codebase. In contrast, our API testing is a system-level testing that emphasizes the dependencies between requests and targets. These dependencies span across functions and files, thereby constituting a repository-level concern.
Repository-level tasks are challenging as information spans across massive code. State-of-the-art works, including package migration  (Bairi et al., 2023), issue resolution  (Ma et al., 2024; Xia et al., 2024; Zhang et al., 2024), and code completion  (Phan et al., 2024; Ding et al., 2023; Shrivastava et al., 2023), employ dependency graphs to identify relevant code components through method/class relationships (e.g., imports, invocations, inheritance). However, different from existing repository-level tasks that focus on capturing module interactions, API testing emphasizes value propagation paths between API requests and testing targets. Specifically, the goal of API testing is to cover the target by modifying the request. Therefore, the key is to determine the target’s expression with respect to the request. Once this expression is established, it provides a constraint representing the target condition. To cover the target, we simply need to satisfy this constraint. To build the expression with respect to the request, we need to track value transmission at the statement level. This analysis works by starting from the variables of the target and iteratively finding their definitions until the variable of the request. Previous method-level or class-level dependency graphs, though useful for identifying related modules, introduce extraneous code elements that obscure the critical value transmission chain. This over-inclusion occurs because class/method dependencies aggregate all interactions within a component, whereas only a subset of these interactions pertain to the specific request-target relationship. Considering the accuracy of semantic representation, as well as the limitations on the length of LLM context and cost considerations, we have opted to conduct code extraction at the statement level.
@@ 3. MioHint
To improve mutation accuracy, we substitute the vanilla mutation with our proposed LLM-assisted mutation. This process involves three main steps, which are related code extraction, prompt construction, and test case generation. Related code extraction retrieves static code related to the target statement from the codebase. Prompt construction combines the information gathered to form a task instruction. Test generation queries LLMs with the constructed prompt, processes the response to get the mutation hint, and applies the mutation hint to the original test case to generate a new one.
In the rest of this section, we first present a high-level overview of MioHint. We then provide a detailed explanation of the related code extraction, prompt construction, and test generation steps.
@@
0:programprogram to test, search time $T$, modelmodel to query for test cases, and minimum count of times to query per iteration $M$.
0: a set of test cases exercising the program that maximizes the coverage.
[…]
4:while $timeElapsed<T$ do
5:  $chosenTarget\leftarrow\text{{ChooseTargetByUncover}}(archive)$
6:  $chosenTestCase\leftarrow\text{{ChooseTestCases}}(testCases,chosenTarget,archive)$
7:  $relatedCode\leftarrow null$
8:  if $\text{{isBranch}}(chosenTarget)\textbf{ or }\text{{isMethodReplacement}}(chosenTarget)$ then
9:   $relatedCode\leftarrow\text{{GetRelatedCode}}(chosenTarget)$
10:  end if
11:  $upToNTimes\leftarrow\text{{GetNumberOfMutations}}()$
12:  $llmTimes\leftarrow 0$
13:  if $relatedCode\neq null$ then
14:   $llmTimes\leftarrow\text{{Max}}(\frac{upToNTimes}{2},M)$
15:  end if
16:  $totalTimes\leftarrow\text{{Max}}(upToNTimes,llmTimes)$
17:  for $i\leftarrow 1$ to totalTimestotalTimes do
18:   if $i\leq llmTimes$ then
19:    $newTestCase\leftarrow\text{{LLMAssistedMutate}}(chosenTestCase,apis,\newline relatedCode,model)$
20:   else
21:    $newTestCase\leftarrow\text{{Mutate}}(chosenTestCase,apis)$
22:   end if
23:   $evaluateResult\leftarrow\text{{calculateCoverage}}(archive,chosenTestCase,\newline newTestCase)$
24:   if $\text{{CoverNotFewer}}(evaluateResult)$ then
25:    $archive\leftarrow\text{{UpdateCoverage}}(archive\cup newTestCase)$
26:    $chosenTestCase\leftarrow\text{{UpdateCurrentTestCases}}(newTestCase)$
27:    $testCases\leftarrow\text{{UpdateTestCases}}(testCases\cup newTestCase)$
28:   end if
29:  end for
30:end while
31:
32:return archivearchive